We ran the cables for the PMC: The RF cable for the 35.5 MHz drive was cheap and so we swapped the 29.5 MHz cable for it.
There now remain 1 RG-174 cable to drive the FSS PC (21.5 MHz) and 3 Heliax for the Kiwamu Tri-Mod EOM (11, 29.5, and 55 MHz).
We also changed the BLACK HV drive cable for the RED one (previously used for the MZ). All HV cables MUST be RED.
The BLACK cable is now used for the PMC_REFL DC.
The Heliax cables are routed onto the table - it remains a Alberto/Kiwamu job to strain relieve them and attach them to the TriMod box and EOM in the morning.
The PMC is locked and we did some partially bootless alignment and mode-matching. It locks easily on a TEM00 mode (with very poor visibility), but the
rest of the beam train can now be aligned while Valera does the PMC matching mambo.